# Mind the Gap: A Readability-Aware Metric for Test Code Complexity

### v1

@@ I Introduction
  • A publicly available dataset of 15,750 generated test suites across 350 Java classes, including CCTR scores, prompt templates, and evaluation scripts: https://anonymous.4open.science/r/CCTR-0A4E/
@@ V-A Empirical Motivation
CCTR is grounded in the developer-centric empirical model proposed by Winkler et al. [13], which identifies three main dimensions of test comprehension:
  • Test structure (e.g., AAA clarity, indentation)
  • Naming and expression semantics (e.g., assertion intent, variable clarity)
  • Test logic and purpose (e.g., relevance of setup, test goal)
Each component of CCTR is designed to reflect one or more of these dimensions. In particular, our inclusion of annotations is inspired by Guerra et al. [26], who found that annotations like @Test and @NotNull improve code readability and convey test roles—but also warned against overuse. CCTR assigns weights accordingly, balancing informativeness with complexity impact.
@@ V-C Evaluation and Observations
Results. Several insights emerge from Table VI:
  • LLM-generated test suites (GPT-4o, Mistral) show moderate complexity (mean 26–30), consistent with their structured yet concise test style.
  • EvoSuite tests score significantly higher, due to their fragmented structure, high method count, and dense assertions.
  • CCTR distinguishes between test suites that SonarSource’s Cognitive Complexity fails to separate—capturing both generator-level and dataset-level differences.
  • Despite variability in LOC and token count, CCTR scales robustly across tools and datasets.
These results support that CCTR better reflects test comprehension effort, aligning with developer-centric principles and structural variation.
@@ V-D Illustration
Each test class represents a different style and level of structural clarity:
  • Nested Trivial Code (Listing LABEL:lst:nested_example): Although semantically trivial, its deep control-flow nesting yields a moderate CCTR score of 12, reflecting structural complexity.
  • LLM-Generated Test Suite (Listing LABEL:lst:llm_test): Composed of 12 concise and semantically clear methods, this test suite also achieves a CCTR score of 12, due to consistent assertions, helpful naming, and informative annotations.
  • EvoSuite-Generated Test Suite (Listing LABEL:lst:evosuite_test): This 17-method class is highly fragmented, with synthetic names and redundant assertions. It reaches a CCTR score of 35, capturing the cognitive effort needed to understand its structure and logic.
These examples demonstrate that CCTR yields differentiated and meaningful scores at the class level, reflecting structural variation across test styles. In combination with results from Table VI, this supports its applicability across both individual examples and large-scale benchmarks.

### v2

@@ I Introduction
  • A publicly available dataset of 15,750 generated test suites across 350 Java classes, including CCTR scores, prompt templates, and evaluation scripts: https://github.com/Cwendkuuni/CCTR
@@ V-A Empirical Motivation
CCTR is grounded in the developer-centric empirical model proposed by Winkler et al. [13], which identifies three main dimensions of test comprehension: test structure (e.g., AAA clarity, indentation), naming and expression semantics (e.g., assertion intent, variable clarity), and test logic and purpose (e.g., relevance of setup, test goal). Each component of CCTR is designed to reflect one or more of these dimensions. In particular, our inclusion of annotations is inspired by Guerra et al. [26], who found that annotations like @Test and @NotNull improve code readability and convey test roles—but also warned against overuse. CCTR assigns weights accordingly, balancing informativeness with complexity impact.
@@ V-C Evaluation and Observations
Results. Several insights emerge from Table VI. LLM-generated test suites (GPT-4o, Mistral) show moderate complexity (mean 26–30), consistent with their structured yet concise test style. In contrast, EvoSuite tests score significantly higher due to their fragmented structure, high method count, and dense assertions. CCTR also distinguishes between test suites that SonarSource’s Cognitive Complexity fails to separate—capturing both generator-level and dataset-level differences. Finally, despite variability in LOC and token count, CCTR scales robustly across tools and datasets. These results support that CCTR better reflects test comprehension effort, aligning with developer-centric principles and structural variation.
@@ V-D Illustration
Each test class represents a distinct style and level of structural clarity. Nested Trivial Code (Listing LABEL:lst:nested_example), while semantically simple, involves deep control-flow nesting and yields a moderate CCTR score of 12, reflecting structural complexity. LLM-Generated Test Suite (Listing LABEL:lst:llm_test) achieves the same score, but for different reasons: concise methods, consistent assertions, clear naming, and informative annotations. In contrast, the EvoSuite-Generated Test Suite (Listing LABEL:lst:evosuite_test) contains 17 fragmented methods with synthetic names and redundant assertions, resulting in a much higher CCTR score of 35, indicative of increased comprehension effort.
These examples show that CCTR produces meaningful, differentiated scores at the class level, reflecting structural variation across test styles. Combined with Table VI, this illustrates its applicability to both individual cases and large-scale benchmarks.